OpenStudy (anonymous):

What are the solutions of 0=x^2+4x-5

OpenStudy (anonymous):

factor

OpenStudy (anonymous):

Is it 5 and -1

OpenStudy (anonymous):

it factors as \[0=(x+5)(x-1)\] yes

OpenStudy (anonymous):

but that means \(x+5=0\implies x=-5\) or \(x-1=0\implies x=1\)

OpenStudy (anonymous):

Ok tks

OpenStudy (anonymous):

in other words the solutions are \(-5\) or \(1\)

OpenStudy (anonymous):

Wait my answer choices are -5 & 1 or 5 & -1

OpenStudy (anonymous):

go with the first one